OpenStudy (anonymous):

Find the break-even point for the given cost and revenue equations. Round to the nearest whole unit. C = 20n + 134,000 R = 160n a. 744 c. 140 b. 180 d. 957

OpenStudy (anonymous):

@AravindG @Blank  @texaschic101

OpenStudy (texaschic101):

I am not too sure how to do this...I think you set them equal to each other and solve for n, which gives you 957. But I am not 100% sure that is how to do it

OpenStudy (blank ):

I got the same. :) 957.142857

OpenStudy (anonymous):

thanks :)

OpenStudy (texaschic101):

20n + 134,000 = 160n -- subtract 20n from both sides 134,000 = 160n - 20n 134,000 = 140n 134,000 /140 = n 957 = n

OpenStudy (anonymous):

Erin has developed a new lotion that will help children with sensitive skin. She invests $15,225 in equipment to manufacture and package her lotion. Each bottle costs $3.25 to manufacture and sells for $12. How many bottles of lotion must she make and sell before her business breaks even? a. 998 bottles b. 1740 bottles c. 174 bottles d. 1269 bottles

OpenStudy (blank ):

Is there a given equation already?

OpenStudy (anonymous):

nope

OpenStudy (blank ):

I got x = 1740. What do the others say?

OpenStudy (blank ):

I put the problem in this equation. 12x - (15225 + 3.25x) = 0

OpenStudy (blank ):

Anyone?

OpenStudy (texaschic101):

15,225 + 325x = 12x 15,225 = 12x - 325x oh...thats gonna be negative

OpenStudy (anonymous):

A local band, The Desert Rats, is planning to make a compact disk. It will cost $12,500 to record and produce a master copy, and an additional $3.00 to make each sale copy. If they plan to sell the final product for $8.00, how many disks must they sell to break even? a. 2500 disks b. 1136 disks c. 1562 disks d. 4166 disks

OpenStudy (texaschic101):

wait...if it costs 3.25 to make each bottle, and they sell it for 12, then they make a profit of (12 - 3.25) = 8.75 per bottle 15,225/8.75 = 1740 (answer is B for the other one)

OpenStudy (blank ):

:)

OpenStudy (blank ):

I was right

OpenStudy (texaschic101):

$3 to make it, $8 to sell it.....they make $5 profit off of every disc 12,500/5 = 2500 discs to break even

OpenStudy (blank ):

Yep. I didn't want to type that all. The answer is A. 2,500 disks.

OpenStudy (texaschic101):

:)

OpenStudy (anonymous):

Sarah is the president of the drama club and wants to publish a newspaper. Expenses are $.90 for printing and mailing each copy, plus $600 total for research and writing. The price of the newsletter is $1.50 per copy. How many copies of the newsletter must the club sell to break even? a. To break even, the drama club must sell 1000 copies. b. To break even, the drama club must sell 100 copies. c. To break even, the drama club must sell 250 copies. d. To break even, the drama club must sell 444 copies.

OpenStudy (blank ):

I got a. To break even, the drama club must sell 1000 copies.

OpenStudy (blank ):

600 + 0.9x = 1.5x x = 1000

OpenStudy (anonymous):

What is the best way to determine whether a given point is a solution to a system of equations? a. substitute the x and y coordinates into one equation and check to see if they work c. substitute the x and y coordinates into both equations and check to see if they work b. graph the equations and see if the point lies on both lines d. solve the system of equations and compare the answer to the given point

OpenStudy (blank ):

I just wondering by c and b are mixed up. :)

OpenStudy (anonymous):

okay I have to go now thanks for all of your help

OpenStudy (anonymous):

ill be on later and will ask for more help from you lol so thank you

OpenStudy (blank ):

answer is c by the way. substitute the x and y coordinates into both equations and check to see if they work